# AlignXIE: Improving Multilingual Information Extraction by Cross-Lingual Alignment

### v1

@@ 5.2 Supervised Evaluation
Task Dataset GPT-4 YAYI-UIE IEPILE AlignXIE
@@
5.3 Zero-Shot Evaluation
Tables 6 and 7 show the zero-shot performance in English and Chinese, respectively. Since our method uses the data and setups of IEPILE, which differ from B2NER in data pruning, exclusion of dataset names, etc., the comparison with B2NER in zero-shot evaluation is unfair. These differences, as shown by Yang et al. 2024, significantly impact zero-shot performance, with the average gaps in English tasks reaching up to $9.3$ points.
AlignXIE achieved SoTA performance across all benchmarks in Chinese tasks, demonstrating substantial and consistent improvements in generalization ability. In English tasks, AlignXIE exhibited a noteworthy average increase of $4.24$ points on the NER task. Similarly, in Chinese tasks, AlignXIE surpassed existing SoTA models on all benchmarks, achieving an average improvement of $3.29$ points. Moreover, the trends demonstrated in both Chinese and English show that the most significant improvement occurs in NER. We hypothesize that the NER alignment training in the first phase is more challenging to generalize to other IE tasks in out-of-domain scenarios, which leaves the introduction of RE and EE alignment training to future work.
Dataset Text-Based Code-Based
ACE 2005 Walker et al. (2006) 84.93 86.07
AnatEM Pyysalo and Ananiadou (2014) 89.63 90.14
BC2GM Kocaman and Talby (2020) 82.53 83.17
BC5CDR Li et al. (2016) 90.03 90.17
CoNLL 2003 Sang and Meulder (2003) 94.67 94.93
WNUT 2017 Derczynski et al. (2017) 66.14 66.03
ResumeNER Zhang and Yang (2018) 94.31 95.37
MSRA Levow (2006) 92.97 94.03
Avg 86.90 87.49
Table 8: Ablation study on code-based method.
5.4 Ablation Study
We conduct comprehensive ablation experiments to investigate whether the code-based multilingual IE and IE cross-lingual alignment phase of AlignXIE contribute to the performance.
Code-Based Multilingual IE
We first evaluate the text-based Wang et al. (2023a) and code-based multilingual IE. Table 8 shows results. To ensure a fair comparison, we remove all class comments in our code-based instructions. We use $6$ English NER datasets, ACE 2005, AnatEM, BC2GM, BC5CDR, CoNLL 2003, WNUT 2017, and $2$ Chinese NER datasets, MSRA, ResumeNER to conduct instruction tuning on the Baichuan2-7B for supervised evaluation. The code-based method surpasses the text-based method by $0.59$ points, with increases of over $1.00$ points in $3$ datasets, especially the Chinese datasets MSRA and ResumeNER. The results show that the code can mitigate schema differences between languages, thereby enhancing the cross-lingual alignment to improve the performance of multilingual IE.
IE Alignment Training
To show whether the IE cross-lingual alignment phase contributes to the performance, we further conduct an ablation study by removing the cross-lingual alignment phase of AlignXIE and denoting it as AlignXIE (w/o align phase). Table 9 shows the average performance of supervised evaluation on $4$ tasks across $45$ benchmarks. The results indicate that, compared to AlignXIE (w/o align phase), AlignXIE achieves a substantial average improvement of $1.1$ points across four tasks. This suggests that the IE cross-lingual alignment phase enhances the overall performance of multilingual IE, demonstrating its contribution to facilitating cross-lingual transfer within the multilingual IE framework.
6 Related Work
Transfer Learning for cross-lingual IE
Huang et al. (2017) introduced a zero-shot learning method for EE, facilitating generalization across unseen languages. Building on zero-shot approaches, Huang et al. (2022) utilized generative language models for cross-lingual EAE. Hsu et al. (2023) enhanced cross-lingual RE through prompt-based learning, reducing reliance on large annotated datasets. Zubillaga et al. (2024) emphasized the significance of typology in cross-lingual EE, particularly for low-resource languages like Basque. In contrast, AlignXIE focuses on leveraging the spontaneous alignment of LLMs to facilitate cross-lingual IE transfer on all IE tasks.
Model NER RE ED EAE Avg
AlignXIE (w/o align phase) 85.06 68.67 74.12 74.00 75.46
AlignXIE 85.63 70.38 75.37 74.87 76.56
Table 9: Ablation study on IE alignment training.
LLMs for Multilingual IE
YAYI-UIE Xiao et al. (2023) proposes a chat-enhanced instruction tuning framework for UIE. IEPILE Gui et al. (2024) collects a comprehensive bilingual IE instruction corpus, and B2NER Yang et al. (2024) further designs a universal entity taxonomy. However, previous works overlook the mutual influence mechanisms between schemas in different languages and lack a unified representation.
Code-Based IE
Code-based IE aims to formulate the IE as a code generation task. CodeIE Li et al. (2023) first leverages the Code-LLM and recasts IE tasks into Python function completion tasks. Code4Struct Wang et al. (2023b) represents schema in Python classes for the EAE task. GoLLIE Sainz et al. (2024) incorporates guidelines within class comments. Besides, KnowCoder Li et al. (2024) utilizes class inheritance, class methods, and type hints to further enrich the schema representation. However, these works focus on the English IE task, and we are dedicated to exploring the significant role of code in multilingual IE.
7 Conclusion
In conclusion, this paper introduces AlignXIE , a code-based multilingual IE model that significantly enhances cross-lingual transfer through unified schema representation and an alignment training phase. AlignXIE utilizes Python classes to represent schemas uniformly across languages, ensuring semantic consistency in multilingual IE tasks. Additionally, our proposed cross-lingual alignment phase leverages high-quality IE parallel datasets ParallelNER, which are generated by our proposed LLM-based pipeline, to boost multilingual generalization. Comprehensive experiments demonstrate that AlignXIE achieves SoTA in Chinese and English IE under various settings, with remarkable cross-lingual capability.
Limitation
Extending multilingual IE training to languages beyond English and Chinese remains an exploration for future work. Meanwhile, the inclusion of additional languages would provide deeper insights into the cross-lingual alignment mechanisms within multilingual IE systems. Introducing parallel data from other IE tasks during the cross-lingual alignment phase also presents promising opportunities for further exploration. Furthermore, due to various limitations, we were unable to retrain the model or conduct a comprehensive evaluation using the B2NER framework. Addressing this issue will be a priority for future work.
Appendix A Experiments
A.1 Pipeline Evaluation
We evaluate our pipeline on the label projection benchmark WikiANN Pan et al. (2017), which is a multilingual NER dataset. We use $10$ languages, including Bengali (bn), German (de), Spanish (es), Persian (fa), Hindi (hi), Korean (ko), Dutch (nl), Russian (ru), Turkish (tr), and Chinese (zh), and randomly sample $1,000$ samples for each language. We compare our pipeline with CLaP Parekh et al. (2024),, Awesome-Align Dou and Neubig (2021), and EasyProject Chen et al. (2023). We mainly evaluate faithfulness following the setting of CLaP, measured as the percentage of instances where translated labels appear in the translated sentence, as accuracy evaluation requires costly native speaker rankings across methods, incurring high costs.
Language Awesome-Align EasyProject CLaP Ours
bn 92 98 93 99
de 99 97 79 99
es 99 99 84 100
fa 96 99 72 100
hi 93 36 90 99
ko 96 93 64 99
nl 99 100 85 100
ru 97 99 66 93
tr 98 98 94 99
zh 92 92 60 99
Avg 96 91 79 99
Table 10: Results of label projection evaluation.
@@ A.1 Pipeline Evaluation
Dataset BERT InsturctUIE IEPILE YAYI-UIE GoLLIE UniversalNER GLiNER KnowCoder GNER B2NER AlignXIE
@@
A.2 Full Comparison of Evaluation
We conduct a comprehensive comparison in English between AlignXIE and all current SoTA methods including monolingual IE methods and multilingual IE methods under the supervised setting in Tables 111213 and zero-shot setting in Table 14. In the NER task, we demonstrate top-$2$ results across $14$ (of $23$) benchmarks; in the RE task, we achieve top-$2$ results across $5$ (of $8$) benchmarks. Moreover, particularly in the ED and EAE tasks, we rank among the top-$2$ results across all benchmarks, further substantiating the significant efficacy of our method in multilingual IE.
Additionally, in the comparison of three benchmarks DIANN Zavala et al. (2018), Ontonotes 5 Hovy et al. (2006), and WNUT 2017 Derczynski et al. (2017), utilized by other English code-based IE work Sainz et al. (2024); Li et al. (2024), AlignXIE outperforms the other baselines, which further demonstrated that cross-lingual alignment can enhance monolingual IE.
In the zero-shot setting, we mainly compare our method with multilingual IE systems including IEPILE, and YAYI-UIE. Due to numerous different settings with B2NER, it is unfair to compare our method with B2NER. Our method achieves superior performance compared to other baseline approaches in the AI subset.
Appendix B Data Statistics
B.1 Dataset Statistics
In this work, we conduct evaluations on $63$ datasets, comprising $34$ datasets for the NER task, $15$ datasets for the RE task, $10$ datasets for the ED task, and $5$ datasets for the EAE task. To fairly compare with code-based methods Knowcoder and GoLLIE in demonstrating the effectiveness of our cross-language alignment, we also used PileNER Zhou et al. (2024) for training. The detailed statics are shown in Tables 15, 16, and 17, respectively.
B.2 ParallelNER
Appendix C Prompts
Description Generation
Using entity description generation as an example, we introduce the prompts we used including Description Initialization and Description Polish in Tables 19 and  20.
Dataset YAYI-UIE IEPILE InsturctUIE KnowCoder AlignXIE
@@ Appendix C Prompts
Task Dataset BERT YAYI-UIE IEPILE KnowCoder AlignXIE
@@
Dataset AI Literature Music Politics Science Avg
GPT4 - - - - - 58.49
YAYI-UIE 52.40 45.99 51.20 51.82 50.53 50.39
IEPILE - - - - - 55.55
B2NER 64.10 69.00 76.40 75.50 76.70 72.34
AlignXIE 65.04 59.68 62.97 63.65 62.29 62.73
InsturctUIE 48.40 48.80 54.40 49.90 49.40 50.18
GoLLIE 60.30 67.10 64.50 60.80 60.50 62.64
UniversalNER 53.50 59.40 65.00 60.80 61.10 59.96
KnowCoder 60.30 61.10 70.00 72.20 59.10 64.54
GliNER 60.60 68.40 69.50 74.80 69.40 68.54
GNER 63.10 68.20 75.70 69.40 69.90 69.26
Table 14: Full comparison of supervised evaluation on Zero-shot NER.
@@
FindVehicle Guan (2022) 21 Traffic 21,547 20,777 20,769
@@
WikiNeural En 3 149,011 92,720 11,590 11,597 115,907
CLUENER2020 Zh 10 19,720 10,000 1,343 1,345 12,688
Table 18: Statistics of ParallelNER datasets.
@@
Table 21: Prompt for Translation.

### v2

@@ Results
Task Dataset GPT-4 YAYI-UIE IEPILE KnowCoder-X
@@
Dataset BERT InsturctUIE IEPILE YAYI-UIE GoLLIE UniversalNER GLiNER KnowCoder GNER B2NER KnowCoder-X
@@
Appendix E Full Comparison of Supervised Evaluation
We conduct a comprehensive supervised comparison in English between KnowCoder-X and all current SoTA methods, including monolingual IE methods and multilingual IE methods under the supervised setting in Tables 1314, and 15. In the NER task, we demonstrate top-$2$ results across $14$ (of $23$) benchmarks; in the RE task, we achieve top-$2$ results across $5$ (of $8$) benchmarks. Moreover, particularly in the ED and EAE tasks, we rank among the top-$2$ results across all benchmarks, further substantiating the significant efficacy of our method in multilingual IE.
Additionally, in the comparison of three benchmarks DIANN Zavala et al. (2018), Ontonotes 5 Hovy et al. (2006), and WNUT 2017 Derczynski et al. (2017), utilized by other English code-based IE work Sainz et al. (2024); Li et al. (2024b), KnowCoder-X outperforms the other baselines, which further demonstrated that cross-lingual alignment can enhance monolingual IE.
Dataset YAYI-UIE IEPILE InsturctUIE KnowCoder KnowCoder-X
@@ Appendix E Full Comparison of Supervised Evaluation
Task Dataset BERT YAYI-UIE IEPILE KnowCoder KnowCoder-X
@@
Appendix F Computational Resource Comparison
In the Universal IE Training Phase, we only train on datasets of the supervised setting, which is the same as the other baselines and necessary for fair comparison. Therefore, the additional computational resources of KnowCoder-X are mainly from the IE Cross-Lingual Alignment Phase. First, this is one of our core contributions, which necessitates additional computational resources. Second, this phase of training is highly efficient. In this phase, we train the model for one epoch with $257$k samples, which is considered a reasonable computational cost. For comparison, the additional computational resources required for the Schema Understanding Training of KnowCoder is one epoch with $926$k samples ($\sim$4x). Meanwhile, KnowCoder even has a large-scale pre-training phase.
Appendix G Data Statistics
G.1 ParallelNER
G.2 Dataset Statistics
In this work, we conduct evaluations on $64$ datasets, comprising $34$ datasets for the NER task, $15$ datasets for the RE task, $10$ datasets for the ED task, and $5$ datasets for the EAE task. To fairly compare with code-based methods, KnowCoder and GoLLIE in demonstrating the effectiveness of our cross-lingual alignment, we also used PileNER Zhou et al. (2024) for training. The detailed statistics are shown in Tables 17, 18, and 19, respectively.
Appendix H Examples
Appendix I Prompts
@@
WikiNeural En 3 149,011 92,720 11,590 11,597 115,907
CLUENER2020 Zh 10 19,720 10,000 1,343 1,345 12,688
Table 16: Statistics of ParallelNER datasets.
@@
FindVehicle Guan et al. (2024) 21 Traffic 21,547 20,777 20,769
@@
Table 22: Prompt for Joint Translation.